The percent by mass of zinc in ZnS is %
Answer #1

ANSWER

Molar mass of ZnS = 65.38 + 32.065 = 97. 445g/mol

Mass percent composition by element, both are in equal number, Zn = 1 atom and S =1 atom.

Mass percent of Zinc in ZnS= 65.38×100/97.445 = 67.094%

Mass percent of Sulphur = 32.065 ×100/97.445 = 32. 906%